OpenStudy (anonymous):

I'm trying to help my son with his math, Fractions no less and well I'm having a hard time understanding the problem. Simple form. The problem is 24/60, I think the simple form is 2/5 but for some reason that answer just doesn't look right to me. Can Some one help.

OpenStudy (mertsj):

Well it is right. good for you

OpenStudy (unklerhaukus):

\[{24\over 60}={3\times 2^3 \over 2^2 \times 3 \times 5}\]\[=2/5\]

OpenStudy (anonymous):

UnkleRhaukus, Thank you. And 2/5 is the final answer correct. It can't go down any more than that right.

OpenStudy (unklerhaukus):